OpenStudy (anonymous):

Use a Double- or Half-Angle Formula to solve the equation in the interval [0, 2π). (Enter your answers as a comma-separated list.)

OpenStudy (anonymous):

sin 2θ + sin θ = 0

OpenStudy (ranga):

sin(2A) = 2sin(A)cos(A) Use this and factor out sine. Equate each factor to zero and solve for theta.

OpenStudy (anonymous):

= 2 sin (theta)cos(theta)

OpenStudy (ranga):

Put it in the original equation, factor out sin(theta)

OpenStudy (anonymous):

I am a little lost that gives me 2sin(theta)=0 cos(theta)=0

OpenStudy (ranga):

Let me use x instead of theta for it is easier to type: sin(2x) + sin(x) = 0 2sin(x)cos(x) + sin(x) = 0 sin(x)(2cos(x) + 1) = 0 sin(x) = 0 or 2cos(x) + 1 = 0 solve for x.

OpenStudy (anonymous):

x=0, or x=1/2

OpenStudy (anonymous):

so sin=0 or cos =1/2

OpenStudy (ranga):

sin(x) = 0 or cos(x) = -1/2 sin(x) = 0 means x = 0, pi (they want x to be in the interval [0, 2pi) cos(x) = -1/2 means x = 2/3pi, 4/3pi So the solutions are: theta = 0, 2/3pi, pi, 4/3pi

OpenStudy (anonymous):

is the pi in 2/3 in the denominator with the three or out to the side?

OpenStudy (ranga):

All pi's are in the numerator.

OpenStudy (anonymous):

I got it correct thanks so much. I do have another question if you would be willing to help me.

OpenStudy (ranga):

\[\Large \Theta = 0, \frac{ 2 }{ 3 } \pi, \pi, \frac{ 4 }{ 3 } \pi\]

OpenStudy (ranga):

I can assist but you should do the work so you can learn.

OpenStudy (anonymous):

alright that is perfectly fine with me. Use an Addition or Subtraction Formula to simplify the equation. sin θ cos 3θ + cos θ sin 3θ = 0

OpenStudy (anonymous):

first I got sin4(theta)= 0 now Im at the second part Find all solutions in the interval [0, 2π). (Enter your answers as a comma-separated list.) and I don't understand how to go from there

OpenStudy (ranga):

Yes, sin(4theta) = 0 is correct. Use your calculator or look up a trig table for sine function and see where all it goes to zero in the interval [0, 2π). If using calculator set it to radians.

OpenStudy (anonymous):

Okay so what actually am I doing?

OpenStudy (anonymous):

its zero at -2pi, -pi, 0, pi, 2pi

OpenStudy (ranga):

find a button labeled arcsin or sin^-1 (sine inverse) amd find the sine inverse for 0.

OpenStudy (ranga):

We are interested only for theta in the interval [0, 2pi) (where 2pi is NOT included because they use a parenthesis ")", but 0 is included in the interval because they use a bracket "["). So 0 and pi are the only values in the interval. So 4(theta) = 0 or pi solve for theta.

OpenStudy (anonymous):

so theta= 0 or pi/4

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

where I type my answers in online for my homework it says im wrong

OpenStudy (ranga):

Did you enter the answer as a comma separated list. That is: 0, pi/4 ?

OpenStudy (anonymous):

yes im thinking there is more then that because in the example that is given it gives a lot of answers. isn't that just the answers for quadrant one don't you have to keep adding until you get through quadrant 4?

OpenStudy (ranga):

Yes that is true because of the 4*theta. So when we look at the arcsin or sine inverse we should not stop at the interval [0, 2pi) but go 4 times as much because of the 4*theta. So look for arcsin(0) from [0, 8pi) first

OpenStudy (ranga):

We should just keep adding pi: So it should be (4theta) = 0, pi, 2pi, 3pi, 4pi, 5pi, 6pi, 7pi so theta = 0, pi/4, pi/2, 3pi/4, pi, 5pi/4, 3pi/2, 7pi/4

OpenStudy (anonymous):

alright perfect and I will definitely be asking you for any more help I need. Thank you very much!

OpenStudy (ranga):

you are welcome.